OpenStudy (anonymous):

WILL GIVE MEDAL, HELP NEEDED. Tara teaches ballet in groups of two students. She rents a room for the ballet classes at $350 per month. She charges $50 per hour per student. Tara teaches five groups of students, and her net earnings per month are $3,150. Part A: If each student learns ballet for the same number of hours, write an equation that can be used to calculate the number of hours each student learns ballet in a month. Part B: How many hours did each student learn ballet each month? Show your work and justify each step of your work.

OpenStudy (anonymous):

Net earnings means the money she keeps after expenses. Since she has only one expense, rent, she has to take in $350 + $3,150 = $3,500. Since she charges $50 per hour, then $50 times the number of hours = $3,500. Dividing $3,500 by $50 = 70 hours. So she has to teach 70 hours at $50 per hour to make $3,500.

OpenStudy (anonymous):

Since each group of students is two and she has five groups of student, the number of students is 2 times 5 = 10 students. Part A: Total number of hours = number of students times hours per student. So the equation looks like this: Let h= hours per student 70=10*h Part B: 70=10*h Given from part A 70/10= (10 * h)/10 Divide both sides by 10 70/10=10/10 * h Commutative property of multiplication 7=1*h simplify 7=h Each student took 7 hours of lessons.